OpenStudy (anonymous):

integrate 4x-7/x^2-x-6 dx

OpenStudy (anonymous):

\[\int\limits_{}^{}4x-7/x ^{2}-x-6 dx\]

OpenStudy (amistre64):

\frac{n}{d} if your doing the latex coding

OpenStudy (amistre64):

(x-3)(x+2) on the denom, then partial it

OpenStudy (amistre64):

it should result in: A ln(x-3) + Bln(x+2) where A and B are the numerators from the partials

OpenStudy (anonymous):

then u set it = to the numerator

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

ok, i'll try that. thank you!

OpenStudy (amistre64):

the "cover up" technique might be simplest

OpenStudy (amistre64):

\[\frac{4x-7}{(x-3)(x+2)}=\frac{A}{x-3}+\frac{B}{x+2}\] \[when\ x=3:\ \frac{4.3-7}{\cancel{(3-3)}(3+2)}=\frac{A=8/5}{x-3}+\frac{B}{x+2}\] \[when\ x=-2:\ \frac{4.-2-7}{(-2-3)\cancel{(-2+2)}}=\frac{8/5}{x-3}+\frac{B=-15/-5}{x+2}\]

OpenStudy (amistre64):

12-7=5 ; not 8 :)

OpenStudy (anonymous):

ok. so far i got B=3

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

and A=1

OpenStudy (amistre64):

A=5/5 = 1 is what i get when i learn to multiply

OpenStudy (anonymous):

ok. i got A=1 too

OpenStudy (anonymous):

i got ln(x-3)+3ln(x+2)

OpenStudy (anonymous):

i had a quick question. how do you know when to use partial fraction decomposition?

OpenStudy (amistre64):

if the top is not the derivative of the bottom, we look to arctan or partials

OpenStudy (anonymous):

oh ok. Thank you :)